Question Help Compute the following probabilities: If Y is distributed N (-2.4). Pr (Ys-5)(Round your response to four decimal places.)
or work (unemployed) in the working-age U.S. population for 2012. Joint Distribution of Employment Status and College Graduation in the U.S. Population Aged 25 and Older, 2012 Unemployed Employed (Y 0) Y1) 0.586 0.639 0.361 1.000 Non-college grads (X = 0) 0.015 Total 0.932 a. Compute E(Y). b. The unemployment rate is the fraction of the labor force that is unemployed. Show that the unemployment rate is given by 1 - E(Y). c. Calculate E(YX 1) and E(Y|X 0) d. Calculate the unemployment rate for () college graduates and (ii) non-college graduates. A randomly selected member of this population reports being unem- ployed. What is the probability that this worker is a college graduate? A non-college graduate? e. f. Are educational achievement and employment status independent? Explain
Answer #1

a) The marginal distribution of Y is given by the column totals in the table

Joint Distribution of Employment Status and College Graduation in the US. Population Aged 25 and Older, 2012 P(x) Unemployed Employed Non-college grads (x- 0) 0.053 College grads (x) Total 0.586 0.346 0.932 Total 0639 0361 1000 0015 Pxy) P(v)0.068

that is P(Y=0)=0.068 and P(Y=1)=0.932

The expected value of Y is

y=0.1 = 0 × 0.068 + 1 × 0.932 0.932

b) From the column totals, we know that if the total labor force is 1, then 0.068 is unemployed. Since unemployment rate is the fraction of the labor force that is unemployed,
Unemployment rate is 0.068

E(Y)=0.932

1-E(Y)=1-0.932 = 0.068

Hence the unemployment rate is given by 1-E(Y)

c) We know the following probabilities

Marginal probabilities of X

P(X=0) = 0.639

P(X=1)=0.361

The joint probabilities

P(X=0,Y=0) = 0.053

P(X=0,Y=1) = 0.586

P(X=1,Y=0) = 0.015

P(X=1,Y=1) = 0.346

The conditional distribution of Y given X=1 is

using the formula for conditional distribution rt-1. У pi r ( 0) _ 0.015-0.0416 0361 = 0.0416 Y = 1) = 0.346 _ 0.9584 P(Y = 0

The conditional expectation of Y given X=1 is

y=0,1 = 0 × 0.0416 1 × 0.9584 0.958

The conditional distribution of Y given X=0 is

using the formula for conditional distribution -o. У = 0) _ 0.053-. 0.0829 0.639 P(X = 0) Pl.X = 0. У-1) P(X=0) 0.586 0.639-0.9171

The conditional expectation of Y given X=0 is

E(Y | X = 0) =〉-yP(y | X = 0) y=0,1 = 0 × 0.0829 + 1 × 0.9171 = 0.917

d). Using the result from part b), we know that unemployment rate is given by 1-E(Y)

i) Unemployment rate for college graduates (X=1) is 1-E(Y|X=1) =1-0.958=0.042

ii) Unemployment rate for non-college graduates (X=0) is 1-E(Y|X=0) =1-0.917=0.083

e) Y=0 is the event that a randomly selected member of this population reports being unemployed.

Let X=1 is the event that a randomly selected member of this population is a college graduate.

Let X=0 is the event that a randomly selected member of this population is a non-college graduate.

the probability that  a randomly selected member of this population is a college graduate (X=1) given that this worker is unemployed (Y=0) is given by

using the formula for conditional distribution 0.015 0.068 0.221

ans: the probability that  a randomly selected member of this population is a college graduate given that this worker is unemployed is 0.221

the probability that  a randomly selected member of this population is a non-college graduate (X=0) given that this worker is unemployed (Y=0) is given by

P(X = 0 | y = 0) = using the formula for conditional distribution P(y-0) 0.053 0.068 0.779

ans: the probability that  a randomly selected member of this population is a non-college graduate given that this worker is unemployed is 0.779

f) 2 events A and B are independent if their joint probability is equal to the product of marginal probability.

That is A and B are independent if

ΡΑλ B) = P(A) × P(B)

We will pick the first cell, the corresponding events are

X=0: selected member of this population is a non-college graduate

Y=0: selected member of this population is a unemployed

The joint probability of X=0 and Y=0 is

P(X = 0. У = 0) = 0.053

The marginal probability of X=0 is P(X=0)=0.639

The marginal probability of Y=0 is P(Y=0)=0.068

the product is

P(X = 0) × P(y = 0) = 0.639 × 0.068 = 0.043

This is not equal to the joint probability.

Hence X=0 and Y=0 are not independent.

We can check this for others, but one is sufficient.

We can say that educational achievement and employment status are not independent.

That is employment status depends on educational achievement.